OpenStudy (anonymous):

what's the value of (x)... 1/2+7x/10=13/20

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{7x}{10}=\frac{13}{20}\] multiply both sides by 20 to get \[10+14x=13\] and then it is much easier

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

thnx